# TetSphere Splatting: Representing High-Quality Geometry with Lagrangian Volumetric Meshes

### v2

@@ Abstract
We present TetSphere splatting, an explicit, Lagrangian representation for reconstructing 3D shapes with high-quality geometry. In contrast to conventional object reconstruction methods which predominantly use Eulerian representations, including both neural implicit (e.g., NeRF, NeuS) and explicit representations (e.g., DMTet), and often struggle with high computational demands and suboptimal mesh quality, TetSphere splatting utilizes an underused but highly effective geometric primitive—tetrahedral meshes. This approach directly yields superior mesh quality without relying on neural networks or post-processing. It deforms multiple initial tetrahedral spheres to accurately reconstruct the 3D shape through a combination of differentiable rendering and geometric energy optimization, resulting in significant computational efficiency. Serving as a robust and versatile geometry representation, TetSphere splatting seamlessly integrates into diverse applications, including single-view 3D reconstruction, image-/text-to-3D content generation. Experimental results demonstrate that TetSphere splatting outperforms existing representations, delivering faster optimization speed, enhanced mesh quality, and reliable preservation of thin structures. Code is available at https://github.com/gmh14/tssplat.
@@ 1 Introduction
Reconstructing 3D geometry stands as a fundamental task in computer vision and graphics. The field has seen significant strides with the advent of diffusion models, showing exceptional capabilities in generating images. This success has spurred research into employing 2D generative methods for 3D reconstruction [67, 46, 94, 10, 80, 50, 53, 49, 56, 68, 47].
These advances prominently featured the use of Eulerian representations, where the geometry is defined on fixed coordinates in the 3D world space. One prevalent category includes neural implicit representations, such as Neural Radiance Fields (NeRF) [58], which uses a density function, and Neural Implicit Surfaces (NeuS) [91], which employs a signed distance function. Both methods are parameterized by neural networks and require an additional mesher step to obtain geometric shapes. These representations take spatial coordinates of the world space as input, inherently classified as Eulerian. In parallel, explicit representation methods such as Deep Marching Tetrahedra (DMTet) [77] and its variants [25, 52] use an explicit deformable tetrahedral grid, with signed distance values defined at the grid vertices – an alternative take on Eulerian representations which facilitates the integration of explicit shading materials. Optimizing Eulerian models, including both implicit and explicit ones, however, is both time- and memory-consuming. This challenge is primarily due to the necessity of employing high-capacity neural networks or high-resolution grids, which are essential for capturing intricate geometric details and inherently contain numerous parameters. Moreover, Eulerian representations are vulnerable in modeling thin structures, often leading to floating artifacts. Extracting surfaces from these models is also required for subsequent applications like rendering and simulation, adding another layer of complexity.
Figure 1: (a) Eulerian vs. Lagrangian geometry representations: DMTet employs a deformable tetrahedral grid, assigning signed distance values at vertices, whereas TetSphere splatting reconstructs 3D shapes by directly deforming tet-spheres, enhancing computational and memory efficiency. TetSphere splatting supports a range of applications, including (b) single-view 3D reconstruction from Google Scanned Objects dataset [19], (c) SDS-based image-to-3D generation, and (d) text-to-3D generation.
To address these challenges, we propose TetSphere splatting, an explicit, Lagrangian geometry representation that efficiently constructs high-quality meshes. Lagrangian representations track the deformation of geometry primitives through space, offering a more efficient approach than the typically more resource-intensive Eulerian representations. An illustrative comparison between Eulerian and Lagrangian geometry representation is shown in Fig. 1 (a). Our TetSphere splatting leverages tetrahedral meshes – a volumetric, Lagrangian representation that remains relatively underexplored in the realm of 3D reconstruction. The 3D object is represented by the union of a set of deformed tetrahedral spheres, termed TetSphere. Compared to Gaussian point cloud, TetSphere imposes structured constraints between points owing to tetrahedralization and also offers a clear distinction between the interior and boundary of the object. Compared to other Lagrangian representations, particularly surface mesh-based methods [1, 65] which struggle with regularization issues leading to self-intersections and compromised mesh quality, TetSphere exhibits greater stability due to its volumetric nature. It also effectively handles complex topologies without re-meshing, an often-necessary requirement in surface mesh representations during optimization [65]. Overall, TetSphere offers the following advantages: 1) its explicit representation ensures fast optimization (in a matter of minutes) and reduced memory cost compared to Eulerian representations; 2) its Lagrangian nature shows robustness to thin structures and reducing floating artifacts; 3) it enables the integration of volumetric regularization, significantly improving mesh quality; and 4) it can handle shapes with arbitrary topologies.
Figure 2: Comparisons of mesh quality across NeRF, NeuS, GS, DMTet, and TetSphere splatting, highlighting a critical yet overlooked criterion of 3D reconstruction. Existing methods are not inherently mesh-based but are derived through post-processing like Marching Cubes. On the contrary, our method demonstrates superior mesh quality without remeshing. Results are obtained from a single image of a dog in Fig. 3.
We further present a computational framework for TetSphere splatting. Similar to Gaussian splatting [41], our method “splats” TetSpheres to conform to the target object. We formulate the deformation of TetSpheres as a geometric energy optimization, incorporating differentiable rendering loss, bi-harmonic energy of the deformation gradient field, and non-inversion constraints, to ensure robust and accurate reconstruction. Our framework is versatile, enabling seamless integration with applications including single-view object reconstruction, image-to-3D generation, and text-to-3D generation. Notably, TetSphere splatting’s fast optimization and reduced computational cost offer significant benefits for the latter two tasks that typically require intensive time and memory.
In our evaluation, we underscore a frequently overlooked aspect of 3D reconstruction: the quality of the reconstructed mesh. The concept of mesh quality encompasses multiple attributes key to the usability of 3D models, such as the uniformity of surface triangles, the absence of undesired bumps, and manifoldness. Despite its paramount importance, particularly in rendering and simulation applications, the evaluation of mesh quality has not received adequate focus in the current 3D reconstruction research. Recognizing this oversight, we introduce three evaluation metrics to assess mesh quality. We conduct a comprehensive evaluation using the Google Scanned Objects (GSO) dataset [19]. Compared to state-of-the-art methods, our TetSphere splatting exhibits superior performance when measured under these newly proposed criteria. It maintains competitive performance on other commonly employed metrics as well. Furthermore, we illustrate its utility in 3D content generation, where it produces qualitatively superior results compared to existing state-of-the-art methods.
@@ 2 Related Work
Eulerian and Lagrangian geometry representations. The differentiation between Eulerian and Lagrangian representations originates from computational fluid dynamics [14] but extends more broadly into computational geometry and physics. Eulerian representations characterize the geometry within a fixed spatial reference, usually utilizing grids or voxels to define the space. In contrast, Lagrangian methods track the movement of individual particles or elements, providing a dynamic representation of geometry that adapts to changes over time. Using fluid simulation as an analogy, an Eulerian view would analyze fluid presence at fixed points in space, whereas a Lagrangian perspective follows specific fluid particles. Neural implicit representations, such as DeepSDF [66], NeRF [59], and InstantNGP [62], are modern adaptations of Eulerian concepts, processing 3D positions as inputs to neural networks. These methods theoretically allow for infinite resolution through NN-based parameterization but can result in slow optimization speeds due to NN training. Beyond implicit ones, explicit or hybrid Eulerian representations, such as DMTet [77], DefTet [24], and TetGAN [102], incorporate explicit irregular grids but can still cause substantial memory usage for high-resolution shapes. Gaussian splatting [82] exemplifies a Lagrangian approach by moving Gaussian point clouds in space. Our proposed TetSphere is an explicit Lagrangian representation, which can be viewed as introducing explicit constraints among points due to tetrahedral meshing, with enhanced efficiency and reconstruction quality.
Single image-to-3D object reconstruction. Single-image 3D reconstruction is an inherently ill-posed problem, and extensive research has been dedicated to addressing it [23, 20]. Early approaches utilized a combination of 2D image encoders and 3D decoders trained on 3D data with both explicit representations, including voxels [13, 88, 12, 97, 98], meshes [90, 26], and point clouds [55, 22, 21, 29], and implicit representations such as NeRF [103, 38, 61], SDF[66, 60, 100], and occupancy networks[57, 5]. Many of these methods were trained on categorized 3D datasets such as 3D templates [72, 27, 40] and semantics [45], yet faced challenges in generalizing to unseen categories. Recently, an active research direction has been leveraging 2D generative models for 3D reconstruction, including the use of SDS and supplementary losses [46, 80, 49, 56, 76, 99, 30, 17]. Alternative approaches train view-conditioned 2D diffusion models to directly produce multi-view images for 3D reconstruction [49, 48, 50, 9, 86, 85, 53]. Our method adopts a similar strategy but introduces an explicit and Lagrangian geometry representation to overcome the limitation of prior representations. The recent introduction of large-scale 3D dataset propelled feed-forward large reconstruction models [33, 92, 101, 96, 81, 75, 15, 16, 8, 41], which directly reconstruct triplane-based NeRF [8] or 3D Gaussians [41]. Their feed-forward inference accelerates the speed of 3D object reconstruction, but often at a sacrifice of relatively low resolution and geometry quality, as well as, coupled geometry and materials.
@@
Our method also relates to text-to-3D content generation as it is one of the applications of TetSphere splatting. We leave a detailed discussion on these related works to Appendix 0.F.
Figure 3: Overall pipeline: TetSphere splatting selects 3D feature points to initialize a set of tetrahedral spheres. The geometry optimization stage deforms these spheres to reconstruct the object. Texture optimization is then applied to obtain the texture or physical material of the object.
3 Overview
We use a set of TetSpheres to represent 3D shapes. Our TetSphere splatting takes multi-view images as inputs. We initiate it using a set of tetrahedral spheres with varying centers and sizes (radii). These spheres are chosen to ensure that, collectively, the initial spheres approximately cover the target shape under different input views. Subsequently, these spheres undergo a two-stage optimization process to precisely reconstruct the 3D object. The first stage, geometry optimization, deforms the tetrahedral spheres through the minimization of rendering loss and two geometric regularization energies. The second stage optimizes the surface texture or Physically Based Rendering (PBR) materials of the tetrahedral spheres. Both stages leverage differentiable rasterizers. The rendering loss typically incorporates supervision from ground truth images or SDS. An illustration of the overall pipeline is shown in Fig. 3.
@@ 4 TetSphere Splatting
Our TetSphere splatting is an explicit, Lagrangian representation for 3D shapes. It positions itself alongside Gaussian splatting [41], which is also a neural-network free Lagrangian representation by representing shapes using Gaussian point clouds as primitives. Although the use of point clouds is notably general, it lacks local constraints within individual Gaussian kernels, resulting in an undefined boundary between the interior and exterior of the object. Converting these Gaussian kernels into explicit meshes necessitates surface reconstruction, which often leads to uneven, bumpy surfaces due to the inherent limitation in defining surface from point clouds.

### v4

@@ Abstract
We introduce TetSphere Splatting, a Lagrangian geometry representation designed for high-quality 3D shape modeling. TetSphere splatting leverages an underused yet powerful geometric primitive – volumetric tetrahedral meshes. It represents 3D shapes by deforming a collection of tetrahedral spheres, with geometric regularizations and constraints that effectively resolve common mesh issues such as irregular triangles, non-manifoldness, and floating artifacts. Experimental results on multi-view and single-view reconstruction highlight TetSphere splatting’s superior mesh quality while maintaining competitive reconstruction accuracy compared to state-of-the-art methods. Additionally, TetSphere splatting demonstrates versatility by seamlessly integrating into generative modeling tasks, such as image-to-3D and text-to-3D generation. Code is available at https://github.com/gmh14/tssplat.
**footnotetext: Both authors contributed equally to this research.
Figure 1: (a) Eulerian vs. Lagrangian geometry representations: Compared to Eulerian methods that rely on a fixed grid, TetSphere splatting, a Lagrangian method, uses a set of volumetric tetrahedral spheres that deform to represent the geometry. TetSphere splatting supports applications such as reconstruction, image-to-3D, and text-to-3D generation (b-d).
@@ 1 Introduction
Accurate 3D shape modeling is critical for many real-world applications. Recent advancements in reconstruction (Mildenhall et al., 2020; Wang et al., 2021b; Kerbl et al., 2023), generative modeling (Poole et al., 2022; Liu et al., 2023b; c; Long et al., 2023), and inverse rendering (Mehta et al., 2022; Nicolet et al., 2021; Palfinger, 2022) have significantly improved the geometric precision and visual quality of 3D shapes, pushing the boundaries of automatic digital asset generation.
Central to these advancements are geometry representations, which can be broadly categorized into two types: Eulerian and Lagrangian representations. Eulerian representations describe a geometry on a set of pre-defined, fixed coordinates in 3D world space, where each coordinate position is associated with properties, such as occupancy within the volume or distance from the surface. Widely used Eulerian representations include neural networks that take continuous spatial coordinates as input to model density fields (Mildenhall et al., 2020) or signed distance functions (Wang et al., 2021b; Mehta et al., 2022), as well as deformable grids that use discrete coordinates, with signed distance values defined at grid vertices (Shen et al., 2021; Gao et al., 2022; Shen et al., 2023b). Despite their popularity, Eulerian representations face a trade-off between computational complexity and geometry quality: Capturing intricate geometric details of the shape requires either a high-capacity neural network or a high-resolution grid, both of which are computationally expensive to optimize in terms of time and memory. This trade-off often limits Eulerian representations when modeling thin, slender structures, as their pre-defined resolution is often insufficient to capture fine details.
Recently, there has been a growing shift in the community towards Lagrangian representations, which are typically more computationally efficient than Eulerian methods (Kerbl et al., 2023; Guédon & Lepetit, 2024; Huang et al., 2024; Son et al., 2024; Chen et al., 2024). Lagrangian representations describe a 3D shape by tracking the movement of a set of geometry primitives in 3D world space. These geometry primitives can be adaptively positioned based on the local geometry of the shape, typically requiring fewer computational resources than Eulerian methods, especially when modeling shapes with fine geometric details. An illustrative comparison between Eulerian and Lagrangian geometry representation is shown in Fig. 1 (a). Two of the most commonly used Lagrangian primitives are 3D Gaussians (Kerbl et al., 2023), which represent geometry using point clouds, and surface triangles (Son et al., 2024; Chen et al., 2024), which are coupled with their connectivity to form surface meshes. While these Lagrangian representations are favored for their computational efficiency, they often struggle with poor mesh quality due to their reliance on tracking individual points or triangles, which can lack overall structural coherence. For instance, Gaussian point clouds can move freely in space, often resulting in noisy meshes, while surface triangles, when coupled with connectivity, can form non-manifold surfaces or irregular, and sometimes degenerated, triangles. The resulting geometry, exhibiting these geometric issues, is unsuitable for downstream tasks such as rendering and simulation, where high-quality meshes are crucial for high fidelity.
Figure 2: Visual comparison of mesh quality across widely used shape representations, including NeRF (Mildenhall et al., 2020), FlexiCubes (Shen et al., 2023b) (Eulerian), DMesh (Son et al., 2024), and Gaussian Splatting (Huang et al., 2024) (Lagrangian). These methods exhibit mesh quality issues, such as irregular or degenerated triangles, non-manifoldness, and floating artifacts. Our method demonstrates uniform surface triangles, improved mesh quality, and structure integrity.
To address these challenges, we propose a novel Lagrangian geometry representation, TetSphere Splatting, designed to construct geometry with an emphasis on producing high-quality meshes. Our key insights stem from the fact that existing Lagrangian primitives are too fine-grained to ensure high-quality meshes. Mesh quality depends not only on individual primitives but also on their interactions: For example, the absence of irregular or degenerated triangles relies on the proper alignment of primitives, while manifoldness depends on how well they are connected. Our representation uses volumetric tetrahedral spheres, termed TetSphere, as geometric primitives. Unlike existing primitives that are individual points or triangles, each TetSphere is a volumetric sphere composed of a set of points connected through tetrahedralization. Initialized as a uniform sphere, each TetSphere can be deformed into complex shapes. Together, a collection of these deformed TetSpheres represents a 3D shape, which is in line with the Lagrangian approach. This more structured primitive allows geometric regularization and constraints to be imposed among points within each TetSphere, ensuring mesh quality is maintained throughout deformation. The volumetric nature of TetSpheres also establishes a cohesive arrangement of points throughout the volume, ensuring structural integrity and effectively reducing common surface mesh issues such as irregular triangles or non-manifoldness.
We further present a computational framework for TetSphere splatting. Similar to Gaussian splatting (Kerbl et al., 2023), our method “splats” TetSpheres to conform to the target shape. We formulate the deformation of TetSphere as a geometric energy optimization problem, consisting of differentiable rendering loss, bi-harmonic energy of the deformation gradient field, and non-inversion constraints, all effectively solvable via gradient descent. For evaluation, we conduct quantitative comparisons on two tasks to assess the geometry quality: multi-view and single-view reconstruction. In addition to the commonly used metrics for evaluating reconstruction accuracy, we introduce three metrics to evaluate mesh quality, focusing on key aspects of 3D model usability: surface triangles uniformity, manifoldness, and structural integrity. Compared to state-of-the-art methods, TetSphere splatting demonstrates superior mesh quality while maintaining competitive performance on other metrics. Furthermore, as demonstrations of its versatility, we showcase TetSphere splatting’s utility in downstream applications of 3D shape generation from both single images and text.
@@ 2 Related Work
Eulerian and Lagrangian geometry representations. The differentiation between Eulerian and Lagrangian representations originates from computational fluid dynamics (Chung, 2002) but extends more broadly into computational geometry and physics. Using fluid simulation as an analogy, an Eulerian view would analyze fluid presence at fixed points in space, whereas a Lagrangian perspective follows specific fluid particles. Neural implicit representations, such as DeepSDF (Park et al., 2019) and NeRF (Mildenhall et al., 2021) are modern adaptations of Eulerian concepts, processing 3D positions as inputs to neural networks. These methods theoretically allow for infinite resolution through NN-based parameterization but can result in slow optimization speeds due to NN training. Explicit or hybrid Eulerian representations, such as DMTet (Shen et al., 2021) and TetGAN (Yang et al., 2019), incorporate explicit irregular grids but can still cause substantial memory usage for high-resolution shapes. Mosaic SDF (Yariv et al., 2024) uses Lagrangian volumetric grids that move in space but are designed for 3D generation tasks only where ground-truth shapes are required. Gaussian splatting (Tang et al., 2023a) exemplifies a Lagrangian approach by moving Gaussian point clouds in space. Surface triangles (Son et al., 2024; Chen et al., 2024) are another example of a Lagrangian approach, where the surface is discretized into a collection of connected triangular elements tracked individually. Our TetSphere can be viewed as introducing constraints among points due to tetrahedral meshing, with enhanced mesh quality.
3D object reconstruction. 3D reconstruction is an inherently ill-posed problem, and extensive research has been dedicated to addressing it (Fu et al., 2021; Fahim et al., 2021). Early approaches utilized a combination of 2D image encoders and 3D decoders trained on 3D data with both explicit representations, including voxels (Chen & Zhang, 2019; Xie et al., 2019; 2020), meshes (Wang et al., 2018; Gkioxari et al., 2019), and point clouds (Mandikal et al., 2018; Groueix et al., 2018), and implicit representations such as NeRF (Yu et al., 2020; Jang & Agapito, 2021; Müller et al., 2022), SDF(Park et al., 2019; Mittal et al., 2022; Xu et al., 2019), and occupancy networks(Mescheder et al., 2019; Bian et al., 2021). Recently, an active research direction has been leveraging 2D generative models for 3D reconstruction, including the use of SDS and supplementary losses (Lin et al., 2023; Sun et al., 2023; Liu et al., 2023b; Melas-Kyriazi et al., 2023; Shen et al., 2023a; Xu et al., 2022; Gu et al., 2023; Deng et al., 2022). The recent introduction of large-scale 3D dataset propelled feed-forward large reconstruction models (Hong et al., 2023b; Wang et al., 2023b; Xu et al., 2023; Weng et al., 2024; Tang et al., 2024; He & Wang, 2023; Tochilkin et al., 2024; Wei et al., 2024; Zhang et al., 2024; Xu et al., 2024b). Their feed-forward inference accelerates the speed of 3D reconstruction, but often at a sacrifice of relatively low resolution and geometry quality. There is also a body of work from the inverse rendering community focused on 3D reconstruction (Nicolet et al., 2021; Palfinger, 2022), including two leveraging explicit Lagrangian representations with differentiable rendering losses (Nicolet et al., 2021; Palfinger, 2022; Vicini et al., 2022; Worchel et al., 2022).  Nicolet et al. (2021) uses Laplacian preconditioning when computing the gradient.  Palfinger (2022) introduces an adaptive remeshing algorithm based on estimating the optimal local edge length. Both methods use a single surface sphere, whereas our method uses multiple tetrahedral spheres as primitives. Detailed discussion and experimental comparisons are provided in Appendix A.
@@
Our method also relates to text-to-3D content generation as it is one of the applications of TetSphere splatting. We leave a detailed discussion on these related works in Appendix J.
Figure 3: Overall pipeline: TetSphere splatting represents a 3D shape using a collection of TetSpheres. Each TetSphere is a tetrahedral sphere that can be deformed from its initial uniform state through deformation gradient. The deformation process is optimized by minimizing rendering loss and geometric energy terms.
@@ 3 TetSphere Splatting
We use tetrahedral spheres as our primitive of choice. Unlike point clouds, tetrahedral meshes enforce structured local connectivity between points owing to tetrahedralization. This preserves the geometric integrity of the 3D shape and also enhances the surface quality by imposing geometric regularization across the entire mesh interior. We formulate the reconstruction of shapes through TetSphere splatting as a deformation of tetrahedron spheres. Starting from a set of tetrahedral spheres, we adjust the positions of their vertices to align the rendered images of these meshes with the corresponding target multi-view images. Vertex movement is constrained by two geometric regularizations on the tetrahedral meshes, derived from the field of geometry processing (Bærentzen et al., 2012). These regularizations, which penalize the non-smooth deformation (via bi-harmonic energy) and prevent the inversion of mesh elements (via local injectivity), have proven effective in ensuring that the resulting tetrahedral meshes are of superior quality and maintain structural integrity. Fig. 3 illustrates the overall pipeline.
3.1 Tetrahedral Sphere Primitive
The primitive of TetSphere splatting is a tetrahedralized sphere, called TetSphere, with $N$ vertices and $T$ tetrahedra. By applying principles from the Finite Element Method (FEM) (Sifakis & Barbic, 2012), the mesh of each sphere is composed of tetrahedral elements, with each tetrahedron constituting a 3D discrete piecewise linear volumetric entity. We denote the position vector of all vertices of the $i$-th deformed sphere mesh as ${x}_{i}\in\mathbb{R}^{3N}$. The deformation gradient of the $j$-th tetrahedron in the $i$-th sphere is denoted as $\mathbf{F}^{(i,j)}_{\mathbf{x}}\in\mathbb{R}^{3\times 3}$, which quantitatively describes how each tetrahedron’s shape transforms (Sifakis & Barbic, 2012). Essentially, the deformation gradient $\mathbf{F}^{(i,j)}_{\mathbf{x}}$ serves as a measure of the spatial changes a tetrahedron undergoes from its original configuration to its deformed state. Refer to Fig. 3 for a visual explanation and Appendix F for an in-depth derivation.
Rather than using a single sphere, our representation employs a collection of spheres to accurately represent arbitrary shapes. Consequently, the complete shape is the union of all spheres. By adopting multiple spheres, this approach ensures that each local region of a shape is detailed independently, enabling a highly accurate representation. Moreover, it allows for the representation of shapes with arbitrary topologies. Such a claim is theoretically guaranteed by the paracompactness property of manifold shapes (James, 2000).
  • Compared to neural representations (e.g., NeRF), our tetrahedral representation does not rely on neural networks, thus inherently accelerating the optimization process.
  • Compared to Eulerian representations (such as DMTet), our approach entirely avoids the need for iso-surface extraction – an operation that often degrades mesh quality owing to the predetermined resolution of the grid space.
  • Compared to other Lagrangian representations, such as Gaussian point clouds and triangle meshes, our method offers a volumetric representation through the use of tetrahedral meshes. Each tetrahedron imposes constraints among vertices, leading to superior mesh quality.
3.2 TetSphere Splatting as Shape Deformation
To reconstruct the geometry of the target object, we deform the initial TetSpheres by changing their vertex positions. Two primary goals govern this process: ensuring the deformed TetSpheres align with the input multi-view images and maintaining high mesh quality that adheres to necessary geometry constraints. Fig. 4 illustrates the iterative process of TetSphere splatting.
To maintain the mesh quality, we leverage bi-harmonic energy – defined in the literature on geometry processing (Botsch & Sorkine, 2007) as an energy quantifying smoothness throughout a field – to the deformation gradient field. This geometric regularization ensures the smoothness of the deformation gradient field across the deformation process, thus preventing irregular mesh or bumpy surfaces. It is important to highlight that this bi-harmonic regularization does not lead to over-smoothness of the final result. This is because the energy targets the deformation gradient field, which measures the relative changes in vertex positions, rather than the absolute positions themselves. Such an approach allows for the preservation of sharp local geometric details, akin to techniques used in physical simulations (Wen et al., 2023). Furthermore, we introduce a geometric constraint to guarantee local injectivity in all deformed elements (Schüller et al., 2013). This ensures that the elements maintain their orientation during the deformation, avoiding inversions or inside-out configurations. This constraint can be mathematically expressed as $\mathrm{det}(\mathbf{F}^{(i,j)}_{\mathbf{x}})>0.$ Importantly, these two terms – bi-harmonic energy for smoothness and local injectivity for element orientation – are universally applicable to any tetrahedral meshes, stemming from their fundamental basis in geometry processing (Bærentzen et al., 2012). More details are discussed in Appendix L.